Find b2 – 4ac and the number of real solutions for the given equation.
16-24+9x^2=0

Answer by stanbon(75887)

Find b2 – 4ac and the number of real solutions for the given equation.
16-24x+9x^2=0
---
a = 9 ; b = -24 ; c = 16
------
b^2 - 4ac = 24^2 - 4*9*16 = 0
----
Since b^2-4ac = 0 there are 2 equal Real Number solutions.
========================
Cheers,
Stan H.
=======================